OpenStudy (anonymous):

Can someone please help me with this problem???? A rectangle table is three times as it is wide. If the area is 75ft^2, find the length and width of the table??

OpenStudy (anonymous):

From the information given you can get the following equations\[L=3W\] \[L \times W=75\]

OpenStudy (anonymous):

Sry, L=3W L*W=75

OpenStudy (anonymous):

then just use substitution to solve for L and W

OpenStudy (anonymous):

3W^2=75

OpenStudy (anonymous):

so W^2 = 25 and W = 5

OpenStudy (anonymous):

let length be 3x and width be x l*w = 75 3x*x = 75 3x^2 = 75 x^2 = 75/3 x = root25 x = 5 so, width = 5 length = 3(5) = 15